At a distance of 3.8 m from a siren of an ambulance, the sound intensity is 3.6 × 10-2 W∙m-2 . Assuming that the siren radiates sound uniformly in all directions, find the total power radiated.


1
Expert's answer
2021-06-23T13:02:50-0400

"I=\frac{P}{4\pi r^2}\to P=4\pi r^2\cdot I=4\cdot3.14\cdot3.8^2\cdot3.6\cdot10^{-2}=6.53\ (W)" . Answer